Plumbers are often paid either by hour or on a per job basis. The rate they receive can affect their working speed. Suppose, they can work either for either $15 per hour, or for $560 plus $3 for each hour less than 40 if they can complete the job in less than 40 hours. Also suppose that it takes T hours to complete the job. Clearly the hourly rate is better if T>=40. If T<40, for what values of T is the hourly rate the better pay scale?
Answer by addingup(3677)

15t = 560+3(40-t)
15t = 560+120-3t
18t = 680
t = 37.77 anything over 37.77: hourly rate is best. If they are asking you to round numbers, at 38 hours hourly is best.
---------------------------
Check:
15(37.7777) = 560+3(40-37.7777)
566.67 = 560+6.67 Correct
:
J